Downton Company manufactures two products, Abby and Mansion. The company prepares its master budget on the basis of standard costs. The following data are for January:

Standards

Abby

Mansion

Direct materials

3 ounces at $14.50 per ounce

4 ounces at $17.30 per ounce

Direct labor

5 hours at $60.50 per hour

6 hours at $81.00 per hour

Variable overhead (per direct labor-hour)

$48.00

$53.50

Fixed overhead (per month)

$368,068

$399,360

Expected activity (direct labor-hours)

6,680

7,800

Actual results

Direct material (purchased & used)

5,200 ounces at $13.50 per ounce

4,600 ounces at $19.75 per ounce

Direct labor

5,110 hours at $64.00 per hour

7.510 hours at $85.60 per hour

Variable overhead

$279,550

$399,510

Fixed overhead

$337,950

$409,500

Units produced (actual)

1,100 units

1,200 units

Compute the appropriate variances. Indicate which variances you would investigate and why.

Abby

Mansion

Amount

U/F

Amount

U/F

Compute the Direct Materials Variances

the direct materials price variance

the direct materials quantity/efficiency variance

the total variance for direct materials

Compute the Direct Labor Variances

the direct labor rate/price variance

the direct labor efficiency variance

the total variance for direct labor

Compute the Variable Manufacturing Overhead Variances

Price variance

Efficiency variance

Compute the fixed manufacturing overhead variances

Price variance

Production volume variance


What I have so far is below, but I'm not sure if I am doing it correctly.

Abby Mansion
Amount U/F Amount U/F
Direct Materials Variances
Direct materials price variance $5,200.00 F ($11,270.00) U
Direct materials quantity/efficiency variance 28,033.33 U (33,446.67) F
Total variance for direct materials ($22,833.33) U $22,176.67 F
Direct Labor Variances
Direct labor rate/price variance $17,885.00 U $34.55 U
Direct labor efficiency variance (23,595.00) F (582,591.69) F
Total variance for direct labor ($5,710.00) F ($582,557.14) F
Variable Manufacturing Overhead Variances
Price variance
Efficiency variance
Fixed Manufacturing Overhead Variances
Price variance
Production volume variance
Answer #1


ADbroer so as per the data and accounting concerts The direct material price variana = (SR-AR) XA Q & Here SR - AR Standard A27550 (0) 3460 (F) NOW Total dired material variana om rate variana + DM quantity variana Abby z - 5200 (F) +22550 (0) 220350Abby Mansion e nis & actud! Output x s.br per unit shis a 1100x5 hr 5500 hrs 1200 x 6 = 7200 hrs (ماذا (7200 - 7510) × 81 (55variable & of efficiency variance. (shrs-A hrs) x SR Mansion Abby S has e actual output xs.hr Per unit z 1200 x 627 2006 a 20variana e e budgeted o/H -abrorbed of 368068-303050 - 65018 (0) - 399360-368640 = 30720(0) fixed 0/H Price variana a budgeted